OpenStudy (anonymous):

can anyone explain and solve this for me ? Write an equation for the height of the point P above the ground as a function of t, time in seconds.????

OpenStudy (anonymous):

 I calculated this so far: the wheel travels 2*pi*5=31.41592654cm in one rotation. If the wheel travels 2*pi*5=31.41592654cm/sec it will take 31.41592654/7=4.488seconds to complete one rotation. The initial height is 5. traveling at 7cm/sec, The height of the center of wheel increase is 2.394736842 Center height equation is 5+2.394736842*t